140 children were watching a magic show. the number of boys were 20 more than twice the number of girls watching the show. how many girls were there?
Answer by algebrahouse.com(1659)

x = number of girls
2x + 20 = number of boys {number of boys was 20 more than twice the number of girls}

x + 2x + 20 = 140 {total number of boys and girls was 140}
3x + 20 = 140 {combined like terms}
3x = 120 {subtracted 20 from each side}
x = 40 {divided each side by 3}

there were 40 girls

For more help from me, visit: www.algebrahouse.com